# Constraints on the trilinear and quartic Higgs couplings from triple Higgs production at the LHC and beyond

### v1

@@ V Conclusions
Our investigation for the prospects at the HL-LHC shows that even though triple-Higgs production is limited by low rates at the LHC, its exploration provides interesting information even if it does not receive additional contributions from new scalar resonances. Bounds can be placed on $\kappa_{4}$ significantly beyond the theoretical constraints from perturbative unitarity.
@@
Acknowledgements — We thank Henning Bahl, Akanksha Bhardwaj, Johannes Braathen, Philipp Gadow, Greg Landsberg and Jürgen Reuter for useful comments and discussions. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2121 “Quantum Universe” - 390833306 and has been partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 491245950.

### v3

@@ V Conclusions
Our investigation of the prospects at the HL-LHC shows that even though triple-Higgs production is limited by low rates at the LHC, its exploration provides interesting information even if it does not receive additional contributions from new scalar resonances. Bounds can be placed on $\kappa_{4}$ significantly beyond the theoretical constraints from perturbative unitarity.
@@
Acknowledgements — We thank Henning Bahl, Akanksha Bhardwaj, Johannes Braathen, Philipp Gadow, Ulrich Haisch, Greg Landsberg, Jürgen Reuter and Tania Robens for useful comments and discussions. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2121 “Quantum Universe” - 390833306 and has been partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 491245950.
Note added: — Shortly after our paper, Ref. Papaefstathiou and Tetlalmatzi-Xolocotzi (2023) appeared on the arXiv which studies triple Higgs production in the $6b$ final state at HL-LHC and FCC, including coupling modifications beyond the Higgs self-couplings.